what is the domain of this rotational function 5X/X-7 ??
Answer by jim_thompson5910(35256)

You cannot divide by zero


So when x = 7, the denominator is 0 (since x-7 = 7-7 = 0)


This means we have to exclude x = 7 from the domain.


So the domain is the set of all real numbers except x cannot equal 7.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim